OpenStudy (anonymous):

How do simply a radical of 50?

OpenStudy (anonymous):

you take out the 2 and it is 2 times radical 25 and 25 goes out to 5 and 5 x 2 is 10 so your answer is 10.

OpenStudy (anonymous):

are you asking this \[\sqrt{50}\]

OpenStudy (anonymous):

Too bad i already failed the question :( Thank you though!

OpenStudy (anonymous):

you're welcome. oh well. :(

OpenStudy (anonymous):

ya. How did you post that like that?

OpenStudy (anonymous):

there is a equation button at down

OpenStudy (anonymous):

Oh, I wish i would of known that earlier! LolThanks.

OpenStudy (anonymous):

oh i wish i knew that too! hahaa

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{25*2}=\sqrt{5^2*2}=5\sqrt{2}\]